OpenStudy (anonymous):

A little trivia. name the code. 1, 1, 2, 3, 5, 8, 13, 21

OpenStudy (anonymous):

Fibonacci

OpenStudy (anonymous):

good good now divide the number 5 by 3 8 by 5 13 by 8 and 21 by 13 what is the irrational number

OpenStudy (loser66):

what is the rational number?

OpenStudy (anonymous):

when u divide the numbers 5 and the previous up u should get around 1.6 and keep doing that through the numbers

OpenStudy (loser66):

don't get the answer. My question is what is the rational number?

OpenStudy (anonymous):

the only rathional numbers are the added 1+1=2+1=2+3=5+3=8

OpenStudy (loser66):

take a look at this http://www.mathsisfun.com/rational-numbers.html

OpenStudy (anonymous):

this is the fibonnaci code 1+1 2 1+2=3 3+2=5 5+3 =8 they are rational except when dividded they usually equal around 1.6

OpenStudy (anonymous):

another code